OpenStudy (anonymous):

The sum of four times a number and six times a second number is 36. The difference of five times the second number and three times the first number is 49. Find the numbers.. Please help!!

OpenStudy (agent0smith):

first number:x second: y sum of four times a number (so... 4x) and six times a second number (sooo 6y) is 36 this means 4x+6y = 36

OpenStudy (agent0smith):

difference of five times the second number (5y) and three times the first number (3x) is 49 difference means subtraction, so 5y - 3x = 49 now you have two equations 4x + 6y = 36 5y - 3x = 49

OpenStudy (wolf1728):

4x + 6y = 36 5y - 3x = 49

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

How do you solve it?

OpenStudy (agent0smith):

4x + 6y = 36 -3x +5y = 49 Using either elimination, or substitution... elimination is easier multiply the whole second equation by 4, and the whole first by 3: 3*4x + 3*6y = 3*36 4*-3x +4*5y = 4*49 now you can add them together...try it (simplify the multiplication first)

OpenStudy (wolf1728):

A) 4x + 6y = 36 B) -3x +5y = 49 Multiply B) by 4/3 -4x + 20/3 y = 65.3333 then Add this to A 4 x + 6y = 36 20/3 y + 6y = 101.3333 12.6666 y = 101.3333 y = 8 4x + 48 = 36 4x = -12 x = -3

OpenStudy (anonymous):

Perfect, thank you!

OpenStudy (wolf1728):

u r welcome